the slope of a line that perpendicular to (2,4)(-1,0)is (A)4/3 (B)-3/4)(C)4/3(D)3/4
Answer by drk(1908)

step 1 - find the original slope. We need to use the slope formula
m = Y2-Y1 / X2-X1.
put your coordinates into the formula, remember coordinates are (X,Y)
m = 0-4 / -1-2
m = -4/-3 = 4/3
step 2 - we want the perpendicular slope. so we take the original slope and apply a negative flip. 4/3 - -> -3/4.
This is answer choice [ B ].